# Omni-Effects: Unified and Spatially Controllable Visual Effects Generation

### v3

@@ Abstract
1 Introduction
@@
Visual effects (VFX) play a crucial role in modern filmmaking, enabling the creation of immersive narratives and fantastical worlds. While traditional VFX pipelines, especially for composite effects requiring simultaneous coordination across different spatial locations, are notoriously complex and resource-intensive Chabanova (2022). Recent rapid advances in video generation technologies Yang et al. (2024); Bao et al. (2024); Kong et al. (2024); Wan et al. (2025) are driving a paradigm shift in VFX creation—transitioning from conventional methods to generative model-powered dynamic, efficient synthesis.
@@
The inherent scarcity of VFX data and pronounced variability in dynamic characteristics across effects pose significant challenges to training generative models. Consequently, current methods Liu et al. (2025b) focus on single-effect generation, employing dedicated Low-Rank Adaptation (LoRA) Hu et al. (2022) tailored to individual effects. However, this paradigm struggles with multi-VFX scenes, exhibiting two critical limitations: Limitation 1. Cross-Adapter Interference where joint multi-LoRA activation Ding et al. (2023) induces spatial occlusion artifacts (Figure 2 (a)) and shared-subspace hybrid training triggers fidelity-degrading cross-effect confusion via task interference Zhang et al. (2025a) (Figure 2 (b, c)). Limitation 2. Spatial-Semantic Misalignment wherein the text-pixel space gap prevents precise spatial cue encoding for VFX placement (Figure 2 (d)). These limitations fundamentally constrain conventional video generation adaptation to complex multiple VFX compositions (multi-VFX).
@@
To address these limitations, we propose Omni-Effects, a unified framework modeling multi-VFX generation as a multi-condition video generation problem, where textual prompts specify effect categories while spatial masks define their precise locations. First, to tackle Limitation 1, we introduce a LoRA-based Mixture of Experts (LoRA-MoE) module Shazeer et al. (2017); Dou et al. (2023); Zhang et al. (2025c), which partitions effects into specialized subspaces, each optimized by dedicated expert branches, with a gating router dynamically activating relevant subspaces to minimize cross-task interference and enhance effect fidelity. Second, to overcome Limitation 2, we present the Spatial-Aware Prompt (SAP), which integrates explicit mask-based spatial conditioning with textual inputs via full-attention mechanisms for accurate effect placement. To mitigate SAP cross-interference during concurrent application, we introduce the Independent-Information Flow (IIF) module, which isolates condition-specific information flows through IIF Attention Mask, preventing unintended effect blending. Collectively, these innovations enable Omni-Effects—to our knowledge, the first VFX framework—to achieve high-fidelity multi-VFX compositions with pixel-level spatial control. (Figure 1).
@@
To advance this research, we construct a high-quality VFX dataset Omni-VFX and develop a specialized VFX data pipeline. This pipeline utilises image editing models Liu et al. (2025a) to generate source image pairs depicting initial/final effect states, which are then synthesised into VFX videos via the FLF2V framework (built on Wan2.1 Wan et al. (2025)). Rigorous manual filtering ensures quality while expanding coverage to distinct effect categories. Further, we introduce an evaluation framework specifically designed for controllable VFX generation tasks. Comprehensive experiments validate the Omni-Effects framework’s superior performance across three core capabilities: single-VFX, multi-VFX, and controllable VFX generation. In summary, the major contributions of our work are as follows:
@@
2 Related Works
2.1 Video Generation Models
Recent advances in diffusion-based video generation Chen et al. (2023); He et al. (2022); Zeng et al. (2024); Kong et al. (2024); Yang et al. (2024); HaCohen et al. (2024); Bao et al. (2024); Polyak et al. (2024); Wan et al. (2025); Teng et al. (2025); Seawead et al. (2025) have enabled text-to-video (T2V) and image-to-video (I2V) synthesis, where input images establish spatial context. As a critical I2V application, VFX generation creates unrealizable fantastical visuals. However, VFX data scarcity forces reliance on Low-Rank Adaptation (LoRA) Hu et al. (2022) for limited-data fine-tuning Liu et al. (2025b). This necessitates separate LoRA models per effect, while combined training causes performance degradation, fundamentally limiting multi-VFX generation within a single video.
@@ 2.1 Video Generation Models
2.2 Conditional Video Generation
Condition-guided diffusion models leverage auxiliary inputs for precise output control, falling into two paradigms: spatial-fusion guidance and high-level semantic guidance. Spatial-fusion methods, exemplified by ControlNet Zhang et al. (2023), integrate condition inputs with denoising inputs. These methods Li et al. (2024); Bai et al. (2024); Bian et al. (2025); Xu et al. (2025b); Lei et al. (2025); Jiang et al. (2025) enable fine-grained spatial alignment while preserving the generation quality of pre-trained diffusion models. High-level semantic methods exploit latent interactions between conditions and the denoising process. Techniques include cross-attention based mechanisms Ye et al. (2023); Zhang et al. (2024); Zhou et al. (2025); Yuan et al. (2025) and conditional token concatenation strategies Wang et al. (2024); Huang et al. (2024b); Tan et al. (2025); Zhang et al. (2025b), dynamically modulate generation through semantic embeddings. Crucially, while these methods effectively handle individual conditions, they struggle to simultaneously and independently control multiple conditions, a critical requirement for professional VFX generation.
@@ 2.2 Conditional Video Generation
3 Method
3.1 Preliminaries
3.1.1 Video Diffusion Models
Video generation models usually utilize the diffusion paradigm Ho et al. (2020); Lipman et al. (2022), which generates samples from a data distribution by progressively denoising samples that are initially drawn from a Gaussian distribution . During training, clean samples undergo iterative corruption through diffusion steps:
@@ 3.1.1 Video Diffusion Models
where is conditions (e.g., text and spatial location), and denotes the condition encoder. By replacing target noise with , which is a weighted combination of and , as the prediction target, the v-prediction Salimans and Ho (2022) is derived, which is adopted in our Omni-Effects framework. Moreover, mainstream video generation leverages Diffusion Transformer (DiT) Peebles and Xie (2023); Ma et al. (2024); Chu et al. (2024) architecture by employing attention mechanisms to model spatiotemporal consistency while aligning conditional inputs with visual outputs Zheng et al. (2024). By integrating diffusion processes with Transformer architectures, the video generation performance is improved, leading to high-quality and accurate video synthesis results.
3.1.2 Spatially Controllable Multi-VFX Generation
@@
where and denote the text and spatial mask encoder, respectively. Notice that, when and the spatial trigger is empty, the above task reduces to the traditional single-VFX generation task Liu et al. (2025b).
3.2 Omni-Effects
To model the above task, video diffusion models require simultaneous support for multi-VFX inference and spatial control capabilities. We accordingly propose Omni-Effects, building upon the CogVideoX Yang et al. (2024) architecture and incorporating two core components: LoRA-MoE and Spatial-Aware Prompt. The overview is illustrated in Figure 3, and the details are as follows.
3.2.1 LoRA-MoE
As mentioned in Figure 2, both multi-LoRA parallel inference and single-LoRA unified training degrade performance. Crucially, we observe the synergistic mechanism in VFX training: compatible VFX-combination training enhances single-VFX generation quality (Figure 4). This discovery motivates our adaptive task-space partitioning strategy: inspired by MoE Shazeer et al. (2017) architectures, we partition distinct effects into specialized subspaces and deploy a gating router for adaptive subspace selection.
Specifically, LoRA-MoE Dou et al. (2023) integrates MoE with LoRA (Figure 3 (a)), which employs an expert ensemble where each LoRA specializes in distinct VFX manifolds. Formally, for input token , a weight is obtained by a gating network for each expert, resulting in , where represents the number of experts. Each expert implements LoRA decomposition:
@@ 3.2.1 LoRA-MoE
This MoE-structured plugin replaces standard FFN linear layers, enabling collaborative expert task-solving. During training, Top- routing strategy () is adopted to enforce exactly non-zero entries in . At inference, all experts are activated to prevent effect suppression caused by Top- filtering, omitting critical experts, which is essential for multi-VFX combination generation. Moreover, to mitigate workload imbalance caused by the gating network favoring a few experts during training, we also employ a balanced routing auxiliary loss Fedus et al. (2022). Comprehensive details are provided in Supplement A. The final training objective is expressed as: , where is hyperparameter.
3.2.2 Spatial-Aware Prompt
For condition , embedding positional descriptors within text prompts proves insufficient for precise spatial control. To investigate this phenomenon, we visualize attention maps across diverse prompts. Crucially, attention consistently activates identical regions regardless of prompt semantics (Figure 5 (a)), evidencing textual position cues’ failure to direct activation toward specified targets. Prior work Liu et al. (2025b); Jiang et al. (2025) mitigates this via ControlNet Zhang et al. (2023) to extract a mask sequence for generation guidance. However, this solution suffers from two critical limitations:
@@ 3.2.2 Spatial-Aware Prompt
4 Data and Training
4.1 Dataset Collection
VFX fundamentally manifest as radical spatio-temporal state transformations (e.g., explosion). Despite modern techniques like animation and Computer Graphics Interface (CGI) Chabanova (2022), modeling such dynamics remains challenging. We introduce a novel pipeline: for any input image, Step1X-Edit Liu et al. (2025a) produces its modified counterpart to establish boundary frames defining a VFX’s initial and terminal states. This constraint provides strong transformation priors for generative models. The FLF2V framework Wan et al. (2025) then synthesizes the final video by compressing VFX production into boundary-constrained state-transition path search, significantly reducing modeling complexity. Through curated manual selection, we build a comprehensive dataset Omni-VFX spanning 55 distinct VFX across instantaneous environmental shifts, artistic styles, human emotions, and so on, enabling diverse creative applications. For more data details, please refer to Supplement B.
4.2 Training
@@
5 Experiments
5.1 Experimental Setup
5.1.1 Evaluation Metrics
Following previous work Liu et al. (2025b), for single-VFX evaluation, we employ two established metrics: Fréchet Video Distance (FVD) Unterthiner et al. (2018) for overall fidelity and Dynamic Degree Huang et al. (2024c) for motion dynamics. For controllable VFX, we introduce three novel metrics. Regional Dynamic Degree (RDD), which utilizes optical flow and masks, quantifies the strength of motion within the target region, thereby quantifies motion strength within target regions to measure visual impression. Effect Occurrence Rate (EOR), which is computed by inputting both the video and a given prompt template into Gemini2.5 Comanici et al. (2025) to obtain the answer, measures intended effect trigger frequency, indicating generation reliability. Building upon EOR, Effect Controllability Rate (ECR) assesses spatial precision by verifying VFX confinement to designated areas. Complete metric details appear in Supplement D.
5.1.2 Implementation Details
Training employs a CogVideoX-5B backbone with LoRA rank of with a total of experts, generating resolution videos. For loss, is set to . We utilize 8 H20 GPUs (96GB) with a batch size of 1 per GPU. We use AdamW Loshchilov and Hutter (2017) at a constant learning rate for steps. During inference, DDIM Nichol and Dhariwal (2021) sampling ( steps) with CFG Ho and Salimans (2022) scale is applied, which can be perform on a single GPU. Extended details are in Supplement C.
5.2 Quantitative Results
5.2.1 Unified VFX Generation
@@ 5.2.2 Controllable VFX Generation
To evaluate our model, we perform comprehensive experiments for single- and multi-VFX control, comparing with state-of-the-art methods including (a) CogVideoX Yang et al. (2024), (b) LTX-Video HaCohen et al. (2024), (c) Wan2.1 Wan et al. (2025), and (d) CogVideoX integrated with ControlNet (CogV+CN). Evaluation targets three spatially localized VFX types—Explode, Melt, and Levitate—to ensure contamination-free assessment.
@@
5.2.3 Generalization
Despite being trained with only effects, our model generalizes to diverse mask conditions during inference using the shared Spatial-Condition LoRA, thereby extending to the generation of more concurrent control VFX (). Omni-Effects demonstrates robust extensibility, successfully handling complex effect combinations (Figures 1 (d) and 8), validating its test-time scalable VFX control capability.
5.3 Ablation Studies
5.3.1 LoRA-MoE
@@ Appendix A Method
A.1 Balanced Routing Auxiliary Loss
Drawing inspiration from Switch Transformers Fedus et al. [2022], we integrate a balanced routing auxiliary loss into LoRA MoE training. Specifically, for a batch with tokens, define as the fraction of tokens routed to expert :
@@ B.1 Dataset Collection
Our dataset construction pipeline, outlined in Figure 9, follows a multi-stage generative approach. Firstly, we define target VFX categories spanning diverse styles (e.g., seasonal transformations, claymation, 3D doll rendering). An initial semantic analysis using Qwen Bai et al. [2023] classifies input images by content and style. Secondly, for each category, Step1X-Edit Liu et al. [2025a] generates stylistically consistent frame pairs (initial-final) conditioned on dynamically constructed prompts. These pairs are then analyzed by Qwen to produce descriptive text prompts, which undergo iterative refinement through Wan2.1-14B’s Video Prompt Extender Wan et al. [2025] for temporal coherence. Thirdly, the optimized prompts and frames jointly drive Wan2.1-14B to synthesize augmented video sequences. Through this pipeline, we achieve coverage of 55 effect categories while maintaining quality via: (1) automated style-consistency checks, and (2) manual validation of visual fidelity.
B.2 Dataset Visualization

### v4

# Omni-Effects: Unified and Spatially-Controllable Visual Effects Generation

@@ Abstract
Code — https://amap-ml.github.io/Omni-Effects.github.io
Extended version — https://arxiv.org/abs/2508.07981
Introduction
@@
Visual effects (VFX) play a crucial role in modern filmmaking, enabling the creation of immersive narratives and fantastical worlds. While traditional VFX pipelines, especially for composite effects requiring simultaneous coordination across different spatial locations, are notoriously complex and resource-intensive (chabanova2022vfx). Recent rapid advances in video generation technologies (Yang2024CogVideoXTD; bao2024vidu; kong2024hunyuanvideo; wan2025) are driving a paradigm shift in VFX creation—transitioning from conventional methods to generative model-powered dynamic, efficient synthesis.
@@
The inherent scarcity of VFX data and pronounced variability in dynamic characteristics across effects pose significant challenges to training generative models. Consequently, current methods (liu2025vfx) focus on single-effect generation, employing dedicated Low-Rank Adaptation (LoRA) (hu2022lora) tailored to individual effects. However, this paradigm struggles with multi-VFX scenes, exhibiting two critical limitations: Limitation 1. Cross-Adapter Interference where joint multi-LoRA activation (ding2023parameter) induces spatial occlusion artifacts (Figure 2 (a)) and shared-subspace hybrid training triggers fidelity-degrading cross-effect confusion via task interference (zhang2025lori) (Figure 2 (b, c)). Limitation 2. Spatial-Semantic Misalignment wherein the text-pixel space gap prevents precise spatial cue encoding for VFX placement (Figure 2 (d)). These limitations fundamentally constrain conventional video generation adaptation to complex multiple VFX compositions (multi-VFX).
@@
To address these limitations, we propose Omni-Effects, a unified framework modeling multi-VFX generation as a multi-condition video generation problem, where textual prompts specify effect categories while spatial masks define their precise locations. First, to tackle Limitation 1, we introduce a LoRA-based Mixture of Experts (LoRA-MoE) module (shazeer2017outrageously; dou2023loramoe; zhang2025context), which partitions effects into specialized subspaces, each optimized by dedicated expert branches, with a gating router dynamically activating relevant subspaces to minimize cross-task interference and enhance effect fidelity. Second, to overcome Limitation 2, we present the Spatial-Aware Prompt (SAP), which integrates explicit mask-based spatial conditioning with textual inputs via full-attention mechanisms for accurate effect placement. To mitigate SAP cross-interference during concurrent application, we introduce the Independent-Information Flow (IIF) module, which isolates condition-specific information flows through IIF Attention Mask, preventing unintended effect blending. Collectively, these innovations enable Omni-Effects—to our knowledge, the first VFX framework—to achieve high-fidelity multi-VFX compositions with pixel-level spatial control. (Figure 1).
@@
To advance this research, we construct a high-quality VFX dataset Omni-VFX and develop a specialized VFX data pipeline. This pipeline utilises image editing models (liu2025step1x) to generate source image pairs depicting initial/final effect states, which are then synthesised into VFX videos via the FLF2V framework (built on Wan2.1 (wan2025)). Rigorous manual filtering ensures quality while expanding coverage to distinct effect categories. Further, we introduce an evaluation framework specifically designed for controllable VFX generation tasks. Comprehensive experiments validate the Omni-Effects framework’s superior performance across three core capabilities: single-VFX, multi-VFX, and controllable VFX generation. In summary, the major contributions of our work are as follows:
@@
Related Works
Video Generation Models
Recent advances in diffusion-based video generation (chen2023videocrafter1; he2022latent; zeng2024make; kong2024hunyuanvideo; Yang2024CogVideoXTD; hacohen2024ltx; bao2024vidu; polyak2024movie; wan2025; teng2025magi; seawead2025seaweed; ling2025vmbench; chen2025s) have enabled text-to-video (T2V) and image-to-video (I2V) synthesis, where input images establish spatial context. As a critical I2V application, VFX generation creates unrealizable fantastical visuals. However, VFX data scarcity forces reliance on Low-Rank Adaptation (LoRA) (hu2022lora) for limited-data fine-tuning (liu2025vfx). This necessitates separate LoRA models per effect, while combined training causes performance degradation, fundamentally limiting multi-VFX generation within a single video.
@@
Conditional Video Generation
Condition-guided diffusion models leverage auxiliary inputs for precise output control, falling into two paradigms: spatial-fusion guidance and high-level semantic guidance. Spatial-fusion methods, exemplified by ControlNet (Zhang2023AddingCC), integrate condition inputs with denoising inputs. These methods (li2024controlnet++; bai2024syncammaster; bian2025videopainter; xu2025hunyuanportrait; lei2025animateanything; jiang2025vace) enable fine-grained spatial alignment while preserving the generation quality of pre-trained diffusion models. High-level semantic methods exploit latent interactions between conditions and the denoising process. Techniques include cross-attention based mechanisms (ye2023ip; zhang2024ssr; zhou2025light; yuan2025identity) and conditional token concatenation strategies (wang2024boximator; huang2024context; tan2025ominicontrol; tan2025ominicontrol2; zhang2025easycontrol), dynamically modulate generation through semantic embeddings. Crucially, while these methods effectively handle individual conditions, they struggle to simultaneously and independently control multiple conditions, a critical requirement for professional VFX generation.
@@
Method
Preliminaries
Video Diffusion Models
Video generation models usually utilize the diffusion paradigm (ho2020denoising; lipman2022flow), which generates samples from a data distribution by progressively denoising samples that are initially drawn from a Gaussian distribution . During training, clean samples undergo iterative corruption through diffusion steps:
@@
where is conditions (e.g., text and spatial location), and denotes the condition encoder. By replacing target noise with , which is a weighted combination of and , as the prediction target, the v-prediction (salimans2022progressive) is derived, which is adopted in our Omni-Effects framework. Moreover, mainstream video generation leverages Diffusion Transformer (DiT) (peebles2023scalable) architecture by employing attention mechanisms to model spatiotemporal consistency while aligning conditional inputs with visual outputs (zheng2024open). By integrating diffusion processes with Transformer architectures, the video generation performance is improved, leading to high-quality and accurate video synthesis results.
Spatially Controllable Multi-VFX Generation
@@
where and denote the text and spatial mask encoder, respectively. Notice that, when and the spatial trigger is empty, the above task reduces to the traditional single-VFX generation task (liu2025vfx).
Omni-Effects
To model the above task, video diffusion models require simultaneous support for multi-VFX inference and spatial control capabilities. We accordingly propose Omni-Effects, building upon the CogVideoX (Yang2024CogVideoXTD) architecture and incorporating two core components: LoRA-MoE and Spatial-Aware Prompt. The overview is illustrated in Figure 3, and the details are as follows.
LoRA-MoE
As mentioned in Figure 2, both multi-LoRA parallel inference and single-LoRA unified training degrade performance. Crucially, we observe the synergistic mechanism in VFX training: compatible VFX-combination training enhances single-VFX generation quality (Figure 4). This discovery motivates our adaptive task-space partitioning strategy: inspired by MoE (shazeer2017outrageously) architectures, we partition distinct effects into specialized subspaces and deploy a gating router for adaptive subspace selection.
Specifically, LoRA-MoE (dou2023loramoe) integrates MoE with LoRA (Figure 3 (a)), which employs an expert ensemble where each LoRA specializes in distinct VFX manifolds. Formally, for input token , a weight is obtained by a gating network for each expert, resulting in , where represents the number of experts. Each expert implements LoRA decomposition:
@@
This MoE-structured plugin replaces standard FFN linear layers, enabling collaborative expert task-solving. During training, Top- routing strategy () is adopted to enforce exactly non-zero entries in . At inference, all experts are activated to prevent effect suppression caused by Top- filtering, omitting critical experts, which is essential for multi-VFX combination generation. Moreover, to mitigate workload imbalance caused by the gating network favoring a few experts during training, we also employ a balanced routing auxiliary loss (fedus2022switch). Comprehensive details are provided in Supplement A. The final training objective is expressed as: , where is hyperparameter.
Spatial-Aware Prompt
For condition , embedding positional descriptors within text prompts proves insufficient for precise spatial control. To investigate this phenomenon, we visualize attention maps across diverse prompts. Crucially, attention consistently activates identical regions regardless of prompt semantics (Figure 5 (a)), evidencing textual position cues’ failure to direct activation toward specified targets. Prior work (liu2025vfx; jiang2025vace) mitigates this via ControlNet (Zhang2023AddingCC) to extract a mask sequence for generation guidance. However, this solution suffers from two critical limitations:
@@
Data and Training
Dataset Collection
VFX fundamentally manifest as radical spatio-temporal state transformations (e.g., explosion). Despite modern techniques like animation and Computer Graphics Interface (CGI) (chabanova2022vfx), modeling such dynamics remains challenging. We introduce a novel pipeline: for any input image, Step1X-Edit (liu2025step1x) produces its modified counterpart to establish boundary frames defining a VFX’s initial and terminal states. This constraint provides strong transformation priors for generative models. The FLF2V framework (wan2025) then synthesizes the final video by compressing VFX production into boundary-constrained state-transition path search, significantly reducing modeling complexity. Through curated manual selection, we build a comprehensive dataset Omni-VFX spanning 55 distinct VFX across instantaneous environmental shifts, artistic styles, human emotions, and so on, enabling diverse creative applications. For more data details, please refer to Supplement B.
Training
@@
Experiments
Experimental Setup
Evaluation Metrics
Following previous work (liu2025vfx), for single-VFX evaluation, we employ two established metrics: Fréchet Video Distance (FVD) (unterthiner2018towards) for overall fidelity and Dynamic Degree (huang2024vbench) for motion dynamics. For controllable VFX, we introduce three novel metrics. Regional Dynamic Degree (RDD), which utilizes optical flow and masks, quantifies the strength of motion within the target region, thereby quantifies motion strength within target regions to measure visual impression. Effect Occurrence Rate (EOR), which is computed by inputting both the video and a given prompt template into Gemini2.5 (comanici2025gemini) to obtain the answer, measures intended effect trigger frequency, indicating generation reliability. Building upon EOR, Effect Controllability Rate (ECR) assesses spatial precision by verifying VFX confinement to designated areas. Complete metric details appear in Supplement D.
Implementation Details
Training employs a CogVideoX-5B backbone with LoRA rank of with a total of experts, generating resolution videos. For loss, is set to . We utilize 8 H20 GPUs (96GB) with a batch size of 1 per GPU. We use AdamW (loshchilov2017decoupled) at a constant learning rate for steps. During inference, DDIM (nichol2021improved) sampling ( steps) with CFG (ho2022classifier) scale is applied, which can be perform on a single GPU. Extended details are in Supplement C.
Quantitative Results
Unified VFX Generation
Controllable VFX Generation
To evaluate our model, we perform comprehensive experiments for single- and multi-VFX control, comparing with state-of-the-art methods including (a) CogVideoX (Yang2024CogVideoXTD), (b) LTX-Video (hacohen2024ltx), (c) Wan2.1 (wan2025), and (d) CogVideoX integrated with ControlNet (CogV+CN). Evaluation targets three spatially localized VFX types—Explode, Melt, and Levitate—to ensure contamination-free assessment.
@@
Generalization
Despite being trained with only effects, our model generalizes to diverse mask conditions during inference using the shared Spatial-Condition LoRA, thereby extending to the generation of more concurrent control VFX (). Omni-Effects demonstrates robust extensibility, successfully handling complex effect combinations (Figures 1 (d) and 7), validating its test-time scalable VFX control capability.
Ablation Studies
LoRA-MoE
SAP+IIF
@@ Appendix A Method
Balanced Routing Auxiliary Loss
Drawing inspiration from Switch Transformers (fedus2022switch), we integrate a balanced routing auxiliary loss into LoRA MoE training. Specifically, for a batch with tokens, define as the fraction of tokens routed to expert :
@@ Appendix B Dataset
Our dataset construction pipeline, outlined in Figure 9, follows a multi-stage generative approach. Firstly, we define target VFX categories spanning diverse styles (e.g., seasonal transformations, claymation, 3D doll rendering). An initial semantic analysis using Qwen (qwen) classifies input images by content and style. Secondly, for each category, Step1X-Edit (liu2025step1x) generates stylistically consistent frame pairs (initial-final) conditioned on dynamically constructed prompts. These pairs are then analyzed by Qwen to produce descriptive text prompts, which undergo iterative refinement through Wan2.1-14B’s Video Prompt Extender (wan2025) for temporal coherence. Thirdly, the optimized prompts and frames jointly drive Wan2.1-14B to synthesize augmented video sequences. Through this pipeline, we achieve coverage of 55 effect categories while maintaining quality via: (1) automated style-consistency checks, and (2) manual validation of visual fidelity.
Dataset Visualization
@@
The distribution of samples across categories is illustrated in Figure 12, while representative visualizations are provided in Figure 11.